OpenStudy (anonymous):

explain in words why 2|x|<-4

OpenStudy (anonymous):

Isn't the less than sign supposed to be the other way around :o

OpenStudy (anonymous):

yeah this is impossible

OpenStudy (anonymous):

whoops forgot the other half of the question. Explain why it ISN'T possible. sorry

OpenStudy (anonymous):

or why does it have no solution

OpenStudy (anonymous):

absolute value can not be zero or lower

OpenStudy (ikram002p):

it can be zero though

OpenStudy (anonymous):

oh, sorry, well, anyways, absolute value cannot be anything lower than 0

OpenStudy (anonymous):

@tcorris Do you get it now?

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (zzr0ck3r):

\(|x|\ge 0>-2\) \(2|x|\ge 2*0=0>2*-2=-4\implies 2|x|>-4\)

OpenStudy (zzr0ck3r):

now write that in words :)